Ask
your own question, for FREE!
Mathematics
8 Online
A regular deck of cards has 52 cards. There are four suits of 13 cards each. The suits are called clubs (black), diamonds (red), hearts (red), and spades (black). Two cards are drawn at random (without replacement) from a regular deck of cards. What is the probability that the first card is red and the second is red?
Still Need Help?
Join the QuestionCove community and study together with friends!
I read the question to mean 2 independent events. After the first card is drawn there are 51 cards remaining of which 25 are red. So 1/2 * 25/51 = 25/102
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
18 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal